When must the grounded conductor be considered a current carrying conductor for the purposes of ampacity adjustment?


What would be the maximum allowable ampacity of each #900MCM Aluminum conductor (XHHW-2 insulation and connected to a device greater than 100A) in a raceway containing 4 of them, where each was to be considered a current carrying conductor?


What would be the maximum allowable ampacity of each #4/0 copper conductor (RHW insulation and connected to a device greater than 100A) in a raceway containing 4 of them, where each was to be considered a current carrying conductor? The raceway is routed along a rooftop, where the ambient temperature is 33°C and the raceway is 3/8” off the surface of the roof.
Answer #1

SOLUTION:

A.    The grounded conductor must be considered a current carrying conductor for the purposes of ampacity adjustment when:

1. when there are more than three current carrying conductors in a raceway

2. grounded conductors that serve unbalanced loads

3. grounded conductors that serve known non-sinusoidal, harmonics generating loads

Thus in short, a grounded conductor must be considered as a current carrying conductor for ampacity adjustments when connected to a system supplying non-linear loads.

--------------------------------------------

B.    What would be the maximum allowable ampacity of each #900MCM Aluminum conductor (XHHW-2 insulation and connected to a device greater than 100A) in a raceway containing 4 of them, where each was to be considered a current carrying conductor?

1. FIND THE AMPACITY OF WIRES

AMPACITY FOR A 900MCM Aluminum conductor XHHW-2 CONDUCTOR = 480 A

AMPACITY FOR FOUR 900MCM Aluminum conductor XHHW-2 CONDUCTOR = 4* 480 A = 1920 A

2. CHOOSE THE ADJUSTMENT FACTOR BASED ON THE NUMBER OF CONDUCTORS BUNDLED TOGETHER IN THE RACEWAY

HERE WE HAVE 4 CONDUCTORS IN A RACEWAY.

THE ADJUSTMENT FACTOR FOR 4 WIRES IS 0.80

THEREFORE DERATED AMPACITY OF ONE 900MCM Aluminum conductor XHHW-2 CONDUCTOR = 480 * 0.80

                                                                                                      = 384 A

--------------------------------------------------------------------------------------------------------------------------

C. What would be the maximum allowable ampacity of each #4/0 copper conductor (RHW insulation and connected to a device greater than 100A) in a raceway containing 4 of them, where each was to be considered a current carrying conductor? The raceway is routed along a rooftop, where the ambient temperature is 33°C and the raceway is 3/8” off the surface of the roof.

## . DE-RATING FACTOR IS CONSIDERED ON WHEN THE AMBIENT TEMPERATURE GOES ABOVE 30 C. ##

TAKING THAT FOR THE GIVEN SYSTEM TEMOERATURE MAY RISE UPTO 60C.

CASE 1: CONSIDERING BOTH AMBIENT TEMERATURE DERATING AND BUNDLED CONDUCTOR DERATING

1. FIND THE AMPACITY OF WIRES

AMPACITY FOR A 4/0 RHW COPPER CONDUCTOR = 230 A AT 75 C

TOTAL NO. OF WIRE IN RACEWAY = 4

2. CHOOSE THE CORRECTION FACTOR BASED ON THE BASED ON THE AMBIENT TEMPERATURE OF 33 C AND CONDUCTORS RATED TEMPERATURE RATING

THE AMBIENT TEMPERATURE CORRECTION FACTOR IS 0.94

3.CHOOSE THE ADJUSTMENT FACTOR BASED ON THE NUMBER OF CONDUCTORS BUNDLED TOGETHER IN THE RACEWAY

THE ADJUSTMENT FACTOR FOR 12 WIRES IS 0.80

THEREFORE DERATED AMPACITY OF ONE 4/0 RHW COPPER CONDUCTOR = 230 A *0.94 * 0.80 = 172.96 A